OpenStudy (solidicecream):

I will give you a medal if you get this right.... This is also my last question for my homework.

OpenStudy (anonymous):

question?

OpenStudy (solidicecream):

\[5\sqrt{28}+7\sqrt{14}/\sqrt{21}-\sqrt{14}\]

OpenStudy (anonymous):

10square root of 7 + 7squareroot of 6 / 3 - squareroot 14

OpenStudy (solidicecream):

Did you use wolframalpha @melissaaaa ?

OpenStudy (anonymous):

whats that ?

OpenStudy (solidicecream):

https://www.wolframalpha.com/ It's a site that sometimes gives you the correct answer to your question.

OpenStudy (anonymous):

oh thats cool , so why didn't you just use that lol

OpenStudy (solidicecream):

The site sometimes gets you the correct answer. He he he it not giving me the answer I want.

OpenStudy (anonymous):

\[5\sqrt{7*2^{2}}+\frac{7 \sqrt{2}*\sqrt{7} }{ \sqrt{3}*\sqrt{7} }-\sqrt{14}\]= \[5*2\sqrt{7}+\frac{ 7\sqrt{2} }{ \sqrt{3} }-\sqrt{14}\]= \[10\sqrt{7}+\frac{ 7*\sqrt{2}*\sqrt{3} }{ \sqrt{3}*\sqrt{3} }-\sqrt{14}\]= \[10\sqrt{7}+\frac{ 7\sqrt{6} }{3 }-\sqrt{14}\]